### v1

# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

Abstract
Generative AI (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation: a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society. This study examines our students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. Evaluation of more than 250 responses with more than 40% providing detailed qualitative feedback reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts, with accompanying recommendations to guide educational institutions in navigating a future driven by GenAI.
1 Introduction
The rapid advancement of artificial intelligence (AI), particularly generative AI (GenAI), is reshaping the educational landscape. As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education. It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear. As GenAI becomes more embedded in academic and professional contexts, understanding how students perceive its influence on future readiness is essential.
This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern. By analyzing responses, this research provides actionable insights into how the GenAI Generation, students most affected by GenAI’s rapid integration, understand its role in shaping their academic and career paths.
Defining the GenAI Generation
The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [14, 6, 19, 5]. As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [5, 6]. This study introduces the concept of the GenAI Generation to describe individuals most directly impacted by the Generative AI Revolution. The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022. The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
  • Developmental Context: Actively undergoing critical cognitive and social growth while navigating the GenAI Revolution.
  • Navigating Uncertainty: Early adopters, engaging with experimental, unregulated technologies without widely established policies.
  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a rapidly changing environment.
  • Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
Figure 1: Illustrates the relationships between key characteristics, emphasizing the ongoing feedback loop that drives development and adaptation amidst uncertainty.
Just as Google transformed how students accessed and retained information, leading to phenomena dubbed the “Google Effect” [31], and the rise of digital technologies sparked concerns about “digital dementia” [32], the GenAI Revolution represents another critical shift in the educational landscape. These past transformations demonstrate how new technologies can reshape cognitive processes, social behaviors, and academic growth, sometimes with unintended, long-term consequences [30, 7, 21]. The GenAI Generation, defined by its unique characteristics, is similarly at the forefront of this era’s challenges and opportunities. Evaluating this cohort is essential to understand how these tools influence learning and development during critical formative years, ensuring that we can build a foundation of guidance, policies, and research to support them and future generations effectively.
Awareness of GenAI’s Capabilities and Applications
For the purpose of this study, students’ awareness of GenAI means their familiarity with tools such as ChatGPT and their understanding of how these tools can be leveraged in educational and professional contexts. Research has revealed varying levels of awareness globally. One study found widespread awareness among students across 76 countries, with significant usage for academic tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity persist [34]. Another study highlighted differences in engagement with GenAI based on demographic factors, underscoring the need for targeted educational support [13]. Research also shows that while GenAI is valued for idea generation and brainstorming, there are reservations regarding its use in completing assignments, emphasizing the need for ethical policies [4].
The transformative potential of GenAI in education has been widely discussed, particularly in its applications for personalized learning and adaptive tutoring systems, which support higher-order cognitive tasks such as problem-solving and critical analysis [9]. However, challenges like data privacy and instructor resistance remain significant barriers to effective integration. Other studies have explored how personalized learning and language support tools can aid students while also warning of potential over-reliance and its impact on critical thinking [20].
Current research on student awareness of GenAI’s impacts reveals perspectives shaped by demographic and geographic factors. These findings highlight the importance of adopting diverse approaches to educate students about GenAI while providing guidance on data privacy and ethical considerations.
Preparedness for an GenAI-Driven Workforce
For the purpose of this study, student preparedness refers to their confidence in navigating a dynamic, AI-driven job market and their perception of how well their education equips them for this future. Research highlights gaps in AI-specific training within higher education curricula, which leave students feeling under-prepared for the evolving workforce [1, 8]. Ethical and accuracy concerns are also noted as barriers to effective integration, further underscoring the importance of incorporating AI literacy and practical applications into academic programs to enhance student readiness [8]. Additionally, project-based and interdisciplinary learning approaches are recommended to prepare students to critically engage with AI tools and align their academic training with workforce demands [10].
Other studies propose professional development programs for educators to ensure that GenAI integration aligns with institutional goals and emphasizes critical digital literacy, enabling both students and faculty to navigate the complexities of GenAI in professional environments [25]. Personalized learning supported by GenAI is also identified as a potential way to improve academic and employability outcomes, though concerns about ethical use and critical thinking remain [15].
Students’ preparedness highlights the interdisciplinary importance of addressing GenAI in the classroom, ensuring clarity and intentionality in these discussions. Although this research does not examine faculty perspectives, it is critical to recognize that educators also require support and guidance to effectively integrate GenAI discussions and tools into their curricula.
Concerns About GenAI’s Potential Implications
For the purpose of this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact. Current research highlights ethical issues such as data privacy and bias, alongside practical challenges like overreliance and the risk of misinformation. Studies have noted strong reservations about using GenAI for completing assignments without disclosure as well as concerns about the uncritical acceptance of AI outputs [4, 2]. Additionally, research underscores how the misuse of GenAI could undermine academic integrity and originality, emphasizing the need for critical evaluation skills and clear institutional guidelines for responsible use [34, 2]. Balancing the benefits of automation with safeguards against biases and ethical risks remains a pressing concern.
Some researchers critique traditional assessment models, advocating for redesigned evaluations that prioritize originality and critical engagement [23]. Others warn that over-reliance on GenAI could diminish critical thinking and intellectual engagement, urging for its balanced integration into educational practices [20]. While GenAI tools have the potential to enhance advanced cognitive processes, concerns persist regarding their impact on foundational learning skills such as critical thinking and engagement [4, 34, 2, 23, 20].
This contrast highlights the urgent need for clear guidance on student use of GenAI across all educational levels. Strategies must be developed to ensure that these tools are employed ethically while fostering constructive critical engagement and safeguarding cognitive development.
Contributions of This Study
This study contributes to the ongoing dialogue about GenAI’s role in shaping students’ academic and professional paths. Centered on awareness, preparedness, and concern, the survey offers a structured view of the GenAI Generation, a group navigating rapid technological change during key years of education. Focusing on higher education, the analysis shows how students weigh GenAI’s potential to support creativity and critical thinking against concerns about overreliance, ethical risks, and reduced engagement. These insights emphasize the need for thoughtful strategies that support ethical and effective use of GenAI in education.
2 Methods
The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers. Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts. Designed for completion in under five minutes [16], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [28]. Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction. While omitting a neutral option is debated [17], this was an intentional design choice to encourage engagement and reduce bias [18]. The survey was administered across two cohorts, with full question text available in the supplemental materials.
The Test 2023 survey was distributed to Computer Science Master’s students via OMSCS (12,000 members) and MSCS (14,000 members) listservs. Invitations were sent on September 28 and 29, 2023, with responses collected through November 1. The Fall 2023 survey was conducted with support from the Office of Academic Effectiveness, which used stratified random sampling to select 5,655 undergraduates representative of the broader student population. Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error. The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
Methods of Analysis
Survey responses were analyzed using correlation, clustering, and free-response evaluation to capture both quantitative and qualitative insights. These methods supported a systematic exploration of trends and perceptions across diverse student groups.
Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis. Degrees were categorized into three groups: Computer Science, Engineering, and Other, based on thematic alignment with disciplinary focus and methodologies. Multidisciplinary fields, such as Robotics and Computational Science and Engineering, were classified under Engineering and Computer Science, respectively, to maintain consistency. This classification enabled cross-disciplinary comparisons and analysis of AI’s impact on skills by degree category and education level.
For the Fall 2023 cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives. The remaining responses totals 102 for the Test 2023 cohort and 115 for the Fall 2023 cohort
Participation rates varied across groups, with Test 2023 at an estimated 0.5%, and Fall 2023 at 2.5%. These variations may reflect the differences in email dissemination frequency and engagement levels.
Correlation Analysis: To identify relationships between variables, a correlation matrix was computed using the Pearson correlation coefficient, measuring the linear relationship between two variables. The absolute value of 0.5 was the threshold set to identify pairs of features with high collinearity, ensuring that variables with substantial overlap were flagged for further consideration.
Optional Response Analysis: An inductive coding approach was used to analyze free-response data. Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words. N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded. Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
3 Results
The Fall Undergraduates group is divided into three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other, totaling 115 respondents. In contrast, the Test Computer Science Graduates group consists solely of 102 respondents in Computer Science.
Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, with CompSci participants slightly more likely to agree at 21%. Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among Test participants at 100%, followed by CompSci and Engineer respondents at 96% and 93%, respectively. When asked about the limitations of generative AI’s potential applications, agreement levels vary, with Engineers and Test respondents showing slightly higher agreement at 34% and 37%. Generative AI experience is highest among Test respondents (93%) and lowest among Engineers (78%).
Figure 2: Responses on GenAI awareness.
Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Fall Undergraduate Computer Science students have the highest percentage at 82%, followed by Test Computer Science Graduates at 77%. Fall Other Majors report 56%, while Fall Engineering students have the lowest percentage at 38%.
Of those that answered “Yes” to job skills changing, there was a follow up question of how soon they believe the changes will occur. Test Computer Science Graduates had the highest expectation of changes occurring the soonest, within 3 years. Fall Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer. The Fall Undergraduate Engineering students mostly expect changes within 5 or 10 years and just 17% expected changes to occur within 3 years.
Figure 3: Respondent counts by degree and cohort.
Figure 4 covers the remaining three questions on GenAI preparedness. Engineering students reported the highest satisfaction with GenAI discussions at 57% and the greatest confidence that professors considered GenAI at 64%. In contrast, Computer Science students had the lowest satisfaction with 38% and the least confidence in faculty engagement at 35%. Despite this, confidence in curriculum preparation was high across all groups, with Engineering leading at 79%. These results suggest stronger GenAI integration in Engineering curricula, while Computer Science students perceive gaps in discussion and faculty consideration.
Figure 4: Responses on preparedness
Concern: The majority view GenAI’s career impact positively, with Test respondents highest at 81% and Engineering, CompSci, and Other lower with 70%, 64%, and 63% respectively. Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%. Other respondents notably lower at 67%. Even with optimism and excitement about positive career impact and excitement, concern about GenAI’s impact was also prevalent, with Other disciplines 74% and Engineering 75%, expressing the most caution.
Correlations: Correlations with an absolute value of 0.5 or higher were examined to highlight meaningful patterns in how students view GenAI’s impact. Among Fall undergraduates, excitement about GenAI benefits is moderately correlated (0.51) with perceiving a positive career impact, and satisfaction with GenAI discussions is similarly correlated (0.51) with confidence in professors’ engagement. In computer science, students who believe changes will occur within ten years tend to also view GenAI’s applications as limited, with a correlation of 0.50. Students who feel prepared are more likely to expect positive career outcomes, with a stronger correlation of 0.60. For engineering students, satisfaction with discussions correlates even more strongly with confidence in faculty preparedness, at 0.64. In other disciplines, excitement about GenAI benefits is negatively correlated at -0.57 with the belief that its applications are limited, and positively correlated at 0.76 with a belief in GenAI’s positive career impact. Additionally, satisfaction with discussions is moderately negatively correlated with expecting imminent changes at -0.53 and with anticipating rapid shifts in required skills at -0.55.
Figure 5: Topics raised in free response
Free Response: Optional free response rates were relatively high, with 40–50% of undergraduates and test graduate students responding to the institutional recommendation prompt, and 23–25% responding to the additional thoughts prompt. Due to limited responses by discipline, the Fall undergraduate cohort was analyzed as a whole.
The most common recommendation—mentioned by over 50% of both cohorts—was teaching students how to use GenAI tools. About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses. Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses.
Emotional tone varied between cohorts. Over half of undergraduates expressed negative emotions, 29% more than the test graduate group, while test graduates were 20% more likely to express positivity. Additional thoughts covered a broader range of themes. Test graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
These open-ended responses provide rich context to the survey data and are further explored in the Discussion. While the first prompt focused on GenAI’s utility, the second invited any remaining reflections from participants.
4 Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers.
Awareness: the uncertainty in early adopters
While overall awareness of GenAI is high, this awareness does not translate uniformly into understanding or confidence. Non-STEM students exhibited the strongest correlation between excitement about GenAI and its perceived career impact, suggesting an optimistic yet potentially less critically informed perspective. In contrast, engineering students had the highest skepticism about GenAI’s practical applications, while computer science students were most likely to view GenAI as a short-lived trend. These viewpoints may reflect the varied depth of engagement with AI technologies in their respective fields, with technical students balancing enthusiasm with caution about GenAI’s limitations and ethical implications.
These nuances are echoed in free responses, where students shared diverse sentiments:
“because aerospace engineering is pretty specialized and oftentimes requires very difficult/complex work, i am ultimately unsure how genai would be used in my field, if at all. however, i trust it can be a valuable tool in the future. i don’t feel like genai has any negative impacts on my future.” — Undergraduate Engineer
“i think it has limited application in anything with more advanced logic or understanding.” — Undergraduate Engineer
“i think it has limited potential because genai currently cannot distinct whether the information it gives is true or not.” — Undergraduate Engineer
“it’s very uncertain and that’s what makes me slightly afraid. i also don’t understand it very well, which i should work on.” — Undergraduate Computer Science
“as far as i can tell, it’s a fad for almost all possible proposed use cases. until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
As seen in some of these responses, there is skepticism of GenAI achieving beyond what the current potential shows, lending to an experience of technology moving faster beyond students’ grasp. Many students rely on self-directed learning, coursework, and personal experiences to form their perspectives on GenAI. This gap in structured AI education leads to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
In the optional free responses, 50% of the students expressed negative emotions, including frustration, fear, and worry, while 20% conveyed uncertainty, with overlapping sentiments highlighting the complex emotional landscape students navigate:
“i work in security. i am not excited about what’s in store.” — Undergraduate Other
“i’m scared it’s going to make my career not what i want it anymore.” — Undergraduate Other
These comments reflect fear and uncertainty about GenAI’s impact, amplified by a lack of classroom discussion. Without institutional guidance, students are left to navigate these changes on their own during critical educational years. Some expressed dissatisfaction with the current state of awareness in their programs overall:
“even addressing its existence in class would be a step forward.” — Undergraduate Engineer
For a transformative tool like GenAI, students are actively seeking guidance to understand its potential, limits, and ethical implications.
Preparedness: addressing the elephant in the classroom
While students recognize GenAI’s growing influence on their careers, many remain unsure whether their education is preparing them for the changes ahead. Unlike future students who may benefit from GenAI-integrated curricula, the current generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
This preparedness gap is reflected in how different groups perceive GenAI’s impact on job skills. Among computer science students, there is a clear pattern: those who feel prepared are most likely to expect a positive career impact.
In contrast, only 56% of ”Other” students and 38% of engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
Despite these varied views, all cohorts reported relatively high confidence in their curricula—possibly reflecting trust in their degree or Institutions’ reputation. Still, engineering students stand out for their skepticism: they are the least likely to expect GenAI-driven skill changes, report the lowest satisfaction with classroom discussions, and express the least confidence in faculty engagement. This gap signals a need for deeper integration and clearer guidance.
Correlations support these patterns. In engineering, classroom discussions on GenAI are linked to higher confidence in professors, suggesting that open engagement builds trust and a sense of readiness. In computer science, skepticism often aligns with delayed expectations, while students who feel prepared tend to view GenAI as a career asset. In the ”Other” cohort, satisfaction with discussions correlates with the belief that GenAI will impact job skills, suggesting confidence in the status quo or comfort with limited exposure.
The free response data adds depth to these findings. Despite overall confidence in their curriculum, over half of students recommended more courses on GenAI and its practical use. Many also called for stronger emphasis on ”fundamentals,” ”critical thinking,” and ethical considerations:
“I think focusing on the idea of working alongside genai on how to take advantage of it to produce better, faster results, but acknowledge its limitations […] and discussing ethical and other ramifications of using ai in certain situations. — Undergraduate Other
“it is a ”tool” that can be used to aid students’ work when used responsibly, and can be integrated in courses like other software tools, such as matlab, cad softwares, etc. but should not replace the fundamentals of engineering that is already being taught, such as hand calculations and understanding the theory.” — Undergraduate Engineer
“strong implementation in to curriculum and learning tools to learn to work with ai and not treat it as a monster.” — Undergraduate Engineer
“[…] the more adapted we are to it and prepared for its growth/advancement, the better off we’ll be to use it as a tool, rather than as a toy. — Undergraduate Other
While many responses showed a proactive approach to GenAI integration, others expressed skepticism or disappointment:
“[…]i’m disappointed that tech won’t prepare me for the near advent of those tools during my classes.” — Undergraduate Engineer
“as far as i can tell, it’s a fad for almost all possible proposed use cases. until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
“show the downsides and flaws of things like gpt-4, discuss environmental impacts, and don’t use it. genai is built on stolen data sets and harms artists. stop endorsing its use. — Undergraduate Other
The findings reveal a misalignment between perceived career impact and readiness. As GenAI begins to enter technical curricula, students in other disciplines risk being left behind. Addressing this gap will require expanding GenAI education across fields, embedding GenAI literacy into non-STEM programs, and equipping faculty to lead informed, inclusive discussions. These efforts are essential to preparing a more equitable and capable generation for a GenAI-driven future.
Concern: learning at what cost?
Student concerns about GenAI range from ethics and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience. The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%). This likely reflects the integration of GenAI into their curriculum and their confidence using such tools.
“[…] genai may be able to do the bulk of the work when designing something functional and usable, so the focus can now be on designing technology that people like and that benefits them. i think genai can help me do what i like, which would benefit people and society.“ — Graduate Computer Science
“im excited for the changes and technological advances it will bring.“ — Graduate Computer Science
In contrast, positive outlook drops to 64% among undergraduate computer science students, reinforcing how exposure shapes perception.
Among undergraduates, belief in a positive career impact ranges from 63% to 70%, with engineering students on the higher end. Excitement is highest among computer science (75%) and engineering students (77%), while the “Other” cohort remains steady at 67%. Even when students are unsure about GenAI’s direct impact on their careers, many still express enthusiasm about its broader potential. Some specific positive takes included:
“i think less people should have to work for as long and more leisure and volunteering time will be helpful for mental health and balance.” — Undergraduate Engineer
“GenAI is the beginning of the 3rd “industrial” revolution. My life has changed a lot from ChatGPT.” — Undergraduate Other
“i think it is important to emphasize both positives and negatives of genai whenever mentioned. ” — Undergraduate Computer Science
Free responses reflect strong concern, especially in the “additional thoughts” section. Over 50% conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
“i think it will actively damage the quality of work and data produced[…]” — Undergraduate Engineer
“to be honest, i really do not care. i feel like genai is going to obliterate most jobs. it almost seems inevitable.[…]the only thing holding me from hopelessness is (blind) faith that my role as a researcher will not be rendered obsolete.” — Undergraduate Computer Science
“i think we are on a precipice of either ruin or greatness. regardless of how we use it, genai will become a powerful tool and have an impact greater than any singular tool since humans gained control of fire.[…] ”— Undergraduate Other
Some students called for regulation:
“[…]we must ensure that there are precautions before the ai reaches a level where its misuse could pose a legitimate harm towards society and overall human advancement.” — Undergraduate Other
Others expressed concern about government use:
“i think the governments will start using genai to produce propoganda in the coming years (if they haven’t already) to manipulate public opinion. or maybe just politicians with bot accounts blindly supporting them and arguing with dissenters on social media.” — Undergraduate Engineer
And some of the more succinct, poignant responses:
“goodbye middle class, we doomed.” — Undergraduate Computer Science
“i’m scared it’s going to make my career not what i want it anymore.” — Undergraduate Other
Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement. Institutions must actively support both students and faculty as they navigate this rapidly evolving landscape.
5 Conclusion
The GenAI Generation, defined as students in their formative educational years during the rise of GenAI, face a rapidly evolving landscape shaped by both opportunity and uncertainty. While many are excited about GenAI’s potential, they also express concerns around preparedness, ethics, and long-term implications. Findings show that students with greater exposure to GenAI, particularly graduate computer science students, report higher confidence and optimism about their future careers. However, such structured exposure is not consistent across disciplines, revealing an institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift.
The GenAI Generation is distinct in its developmental context as students are not only adapting to new technologies but doing so without established norms or clear guidelines. They are the early adopters of tools like ChatGPT, learning to balance traditional skills with new competencies required to interact with GenAI effectively. Their adaptability is frequently tested while grappling with ethical questions, workforce changes, and broader societal implications.
Students want more than just exposure: they seek guidance on using GenAI tools responsibly, critical thinking around AI-generated outputs, and a robust understanding of ethical considerations. The GenAI Generation is asking for a holistic approach that not only enhances technical skills but also nurtures critical engagement and ethical reasoning. It is critical to closely monitor and derive clarity on the potential impacts of these tools, developing a deeper understanding of cognition and ethical implications.
The GenAI Generation is at a pivotal moment. As the first cohort to learn and grow in a GenAI-centric world, their experiences will set the tone for future generations. The data highlights the need for GenAI literacy efforts that span disciplines and prepare both students and faculty to engage with these tools in meaningful ways. The voices of these students are clear: they want to be part of a future where GenAI is a tool for growth, not a source of uncertainty, and where their education equips them to use GenAI as a force for positive change.
Acknowledgements
We would like to thank the students who took the time to respond to our survey, Dr. Joseph Ludlum and Dr. Loraine Phillips from the Office of Academic Effectiveness at Georgia Tech for their support in distributing the Fall 2023 cohort survey, and Dr. Thomas Ploetz for the continued encouragement throughout the process.