### v2

# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

Abstract
Generative Artificial Intelligence (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation. We define the GenAI Generation as a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society. This study examines students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. Notably, readiness appears increasingly tied to exposure to GenAI through one’s coursework. Students with greater curricular exposure to GenAI tend to feel more prepared, while those without it more often express vulnerability and uncertainty, highlighting a new and growing divide in readiness that goes beyond traditional disciplinary boundaries. Evaluation of more than 250 responses, with over 40% providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts. The challenge ahead involves implementing associated recommendations for educational institutions, moving beyond the baseline of access toward more informed guidance on the use of these tools, while preserving critical thinking, ethical reasoning, and adaptive learning.
Keywords: Generative Artificial Intelligence; GenAI Generation; Student Perceptions; Higher Education; Career Readiness; Cognitive Development; Adaptability; Problem-Solving; Critical Thinking; Ethics in AI
1 Introduction
The significant advancement of Artificial Intelligence (AI), particularly Generative AI (GenAI), is reshaping the educational landscape. GenAI is a type of AI trained on large datasets to generate new content such as text, images, audio, or video in response to user prompts, setting it apart from predictive models by its ability to create [22, 28, 20, 21]. As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education. It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear. As GenAI continues to evolve exponentially and integrate within academic contexts, there is a pressing need to understand how it is shaping the developmental experiences of students in their formative academic years. Unlike previous waves of technology’s impact on education, GenAI’s pace and scale introduce a layer of uncertainty around its long-term effects on cognition, critical thinking, and engagement. By honing in on student perspectives, this research provides early insight into how GenAI is influencing not only perceived career readiness but also the cognitive processes underpinning learning and decision-making.
This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern. By analyzing responses, it provides actionable insights into how the GenAI Generation—students most directly impacted by the rise of GenAI—understand its role in shaping both their academic experiences and future trajectories.
The GenAI Generation
The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [11, 5, 16, 4]. As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [4, 5]. This study introduces the GenAI Generation. The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022. The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
  • Developmental Context: Actively undergoing critical cognitive and social growth while navigating the GenAI Revolution.
  • Navigating Uncertainty: Early adopters, engaging with experimental, unregulated technologies without widely established policies.
  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a highly dynamic environment.
  • Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
Figure 1: Illustrates the relationships between key characteristics, emphasizing the ongoing feedback loop that drives development and adaptation amidst uncertainty.
Just as Google transformed how students accessed and retained information, leading to phenomena dubbed the “Google Effect” [26], and the rise of digital technologies sparked concerns about “digital dementia” [27], the GenAI Revolution represents another critical shift in the educational landscape. These past transformations demonstrate how new technologies can reshape cognitive processes, social behaviors, and academic growth, sometimes with unintended, long-term consequences [25, 6, 18]. The GenAI Generation, defined by its unique characteristics, is similarly at the forefront of this era’s challenges and opportunities. Evaluating this cohort is essential to understand how these tools influence learning and development during critical formative years, ensuring that we can build a foundation of guidance, policies, and research to support them and future generations effectively.
Awareness of GenAI’s Capabilities and Applications
For the purpose of this study, students’ awareness of GenAI means their familiarity with tools such as ChatGPT and their understanding of how these tools can be leveraged in educational and professional contexts. Research has revealed varying levels of awareness globally. One study found widespread awareness among students across 76 countries, with significant usage for academic tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity persist [29]. Another study highlighted differences in engagement with GenAI based on demographic factors, underscoring the need for targeted educational support [10]. Research also shows that while GenAI is valued for idea generation and brainstorming, there are reservations regarding its use in completing assignments, emphasizing the need for ethical policies [3].
The transformative potential of GenAI in education has been widely discussed, particularly in its applications for personalized learning and adaptive tutoring systems, which support higher-order cognitive tasks such as problem-solving and critical analysis [8]. However, challenges like data privacy and instructor resistance remain significant barriers to effective integration. Other studies have explored how personalized learning and language support tools can aid students while also warning of potential over-reliance and its impact on critical thinking [17].
Current research on student awareness of GenAI’s impacts reveals perspectives shaped by demographic and geographic factors. These findings highlight the importance of adopting diverse approaches to educate students about GenAI while providing guidance on data privacy and ethical considerations.
Preparedness for an GenAI-Driven Workforce
For the purpose of this study, student preparedness refers to their confidence in navigating a dynamic, AI-driven job market and their perception of how well their education equips them for this future. Research highlights gaps in AI-specific training within higher education curricula, which leave students feeling under-prepared for the evolving workforce [1, 7]. Ethical and accuracy concerns are also noted as barriers to effective integration, further underscoring the importance of incorporating AI literacy and practical applications into academic programs to enhance student readiness [7]. Additionally, project-based and interdisciplinary learning approaches are recommended to prepare students to critically engage with AI tools and align their academic training with workforce demands [9].
Other studies propose professional development programs for educators to ensure that GenAI integration aligns with institutional goals and emphasizes critical digital literacy, enabling both students and faculty to navigate the complexities of GenAI in professional environments [23]. Personalized learning supported by GenAI is also identified as a potential way to improve academic and employability outcomes, though concerns about ethical use and critical thinking remain [12].
Students’ preparedness highlights the interdisciplinary importance of addressing GenAI in the classroom, ensuring clarity and intentionality in these discussions. Although this research does not examine faculty perspectives, it is critical to recognize that educators also require support and guidance to effectively integrate GenAI discussions and tools into their curricula.
Concerns About GenAI’s Potential Implications
For the purpose of this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact. Current research highlights ethical issues such as data privacy and bias, alongside practical challenges like overreliance and the risk of misinformation. Studies have noted strong reservations about using GenAI for completing assignments without disclosure as well as concerns about the uncritical acceptance of AI outputs [3, 2]. Additionally, research underscores how the misuse of GenAI could undermine academic integrity and originality, emphasizing the need for critical evaluation skills and clear institutional guidelines for responsible use [29, 2]. Balancing the benefits of automation with safeguards against biases and ethical risks remains a pressing concern.
Some researchers critique traditional assessment models, advocating for redesigned evaluations that prioritize originality and critical engagement [19]. Others warn that over-reliance on GenAI could diminish critical thinking and intellectual engagement, urging for its balanced integration into educational practices [17]. While GenAI tools have the potential to enhance advanced cognitive processes, concerns persist regarding their impact on foundational learning skills such as critical thinking and engagement [3, 29, 2, 19, 17].
This contrast highlights the urgent need for clear guidance on student use of GenAI across all educational levels. Strategies must be developed to ensure that these tools are employed ethically while fostering constructive critical engagement and safeguarding cognitive development.
Contributions of This Study
This study offers timely insights into how students in their formative educational years are experiencing the rapid rise of Generative AI. It introduces the term “GenAI Generation” to describe students navigating both academic development and technological transformation without consistent institutional guidance or established norms.
Drawing from a combination of structured survey data and rich qualitative responses, the study reveals meaningful patterns in students’ awareness, preparedness, and concern regarding GenAI. Across majors, students expressed a desire for more intentional support, not just access to GenAI tools, but structured opportunities to learn how to use them responsibly. While formal curriculum integration remains inconsistent, students who encounter GenAI through coursework or self-directed exploration are more likely to articulate confidence and optimism about its role in their future.
From enthusiasm to anxiety, the findings illuminate the psychosocial and pedagogical dimensions of GenAI’s presence in higher education. Students are not only engaging with a new technology, they are forming perceptions of their own readiness, trust in institutions, and ethical responsibility in real time. These perspectives carry clear implications for curriculum design, faculty preparation, and equitable access to digital literacy.
This work contributes to an emerging understanding of what students need to thrive in a GenAI-influenced world: not only technical fluency, but critical engagement, ethical grounding, and institutional clarity. The study underscores the importance of proactive educational responses that extend beyond STEM fields to support students across all disciplines in this pivotal moment.
2 Methods
The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers. Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts. This study received Institutional Review Board (IRB) approval from the [Institution Name Redacted for Review], and all participants provided informed consent prior to beginning the survey. Designed for completion in under five minutes [13], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [24]. Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction. While omitting a neutral option is debated [14], this was an intentional design choice to encourage engagement and reduce bias [15]. The survey was administered across two [Institution Name Redacted for Review] student cohorts, with full question text available in the supplemental materials.
The first cohort, Fall 2023 Computer Science Graduate students, survey was distributed to Online Computer Science Master’s (OMSCS) students, approximately 12,000 members [CITE REDACTED FOR REVIEW], and campus Computer Science Master’s (MSCS), approximately 14,500 members [CITE REDACTED FOR REVIEW], via listservs. Invitations were sent on September 28 and 29, 2023, with responses collected through November 1.
The second cohort, Fall 2023 Undergraduate students, survey was conducted with support from the [Institution Name Redacted for Review] Office of Academic Effectiveness (OAE), which used stratified random sampling to select 5,655 undergraduates representative of the broader student population. Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error. The demographic breakdown of the sample can be found in the supplemental materials. The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
Analysis Approach
Descriptive statistics and frequency distributions were used to compare response patterns across student subgroups, while qualitative evaluation of open-ended responses enabled a systematic exploration of emerging themes and perceptions.
Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis. Where the first cohort of graduate students comprised soley of computer science majors, the undergraduate cohort had many degrees. Rather than broadly distinguishing between STEM and non-STEM, degrees were categorized into Computer Science, Engineering, and Other, based on disciplinary focus and methodological orientation. Table 1 provides a Major-Category mapping for reference. This grouping structure reflects both practical sample sizes (Other, 28 Computer Science, and 60 Engineering) and a meaningful distinction in how students may engage with GenAI in academic contexts. This also enabled a comparison between computer science graduate and undergraduate perspectives.
Grouping a diverse range of degrees—such as Biology, Business Administration, Literature, and Public Policy—under the single category “Other” introduces a potential limitation, as it may obscure important disciplinary differences in students’ exposure to or perspectives on GenAI. While specific exposure to AI tools and curricula was not measured directly, these categories serve as a reasonable proxy for varying degrees of disciplinary integration with GenAI-related content, allowing exploratory comparisons across educational domains.
Table 1: Undergraduate Sample and Response Rates by Majors and Category
Major* Category Sample Freq. Response Freq. Sample % Response %
Applied Physics Other 12 2 0.2 1.7
Biology Other 162 3 2.9 2.6
Business Administration Other 440 3 7.8 2.6
Computational Media Other 98 1 1.7 0.9
Earth & Atmospheric Sciences Other 13 1 0.2 0.9
Econ & Int’l Affairs Other 15 1 0.3 0.9
Economics Other 49 2 0.9 1.7
History, Technology, & Society Other 18 2 0.3 1.7
Industrial Design Other 88 2 1.6 1.7
Int’l Affairs & Mod Lang Other 18 2 0.3 1.7
Lit., Media, & Communication Other 59 2 1.0 1.7
Mathematics Other 77 3 1.4 2.6
Neuroscience Other 146 3 2.6 2.6
Physics Other 68 5 1.2 4.3
Psychology Other 66 5 1.2 4.3
Public Policy Other 37 1 0.7 0.9
Aerospace Engineering Engineer 349 8 6.2 7.0
Biomedical Engineering Engineer 399 6 7.1 5.2
Chemical and Biomolecular Eng Engineer 187 5 3.3 4.3
Computer Engineering Engineer 187 2 3.3 1.7
Electrical Engineering Engineer 187 3 3.3 2.6
Environmental Engineering Engineer 58 3 1.0 2.6
Industrial Engineering Engineer 344 5 6.1 4.3
Materials Science & Engr Engineer 67 3 1.2 2.6
Mechanical Engineering Engineer 582 12 10.3 10.4
Nuclear & Radiological Engr Engineer 32 2 0.6 1.7
Computer Science Computer Science 1419 28 25.1 24.3
Applied Lang/Intercultural St No Response 4 0 0.1 0.0
Architecture No Response 76 0 1.3 0.0
Atmospheric & Oceanic Sciences No Response 3 0 0.1 0.0
Biochemistry No Response 88 0 1.6 0.0
Building Construction No Response 13 0 0.2 0.0
Chemistry No Response 40 0 0.7 0.0
Environmental Science No Response 10 0 0.2 0.0
Global Econ/Mod Lang No Response 5 0 0.1 0.0
International Affairs No Response 22 0 0.4 0.0
Music Technology No Response 23 0 0.4 0.0
Solid Earth/Planetary Sciences No Response 2 0 0.0 0.0
Special/Non-Degree No Response 64 0 1.1 0.0
Civil Engineering No Response 128 0 2.3 0.0
*For reference; sample was not balanced on this attribute. Sample breakdowns by additional attributes can be found within the supplemental material.
For the undergraduate cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives. The remaining responses totals 102 for the graduate cohort and 115 for the undergraduate cohort. Table 2 shows this process
Table 2: Table showing data cleaning process
Survey Population Total Started Total Incomplete Total Dropped Reason for Drop Remaining Completed Total
Computer Science Graduates 2023 26,000** 122 19 1 Response was “prefer not to answer” for degree 102
Undergraduates 2023 5,655 143 26 2 Did not have “undergraduate” for degree type 115
** Approximate total population estimate for Computer Science Graduate 2023 cohort.
Qualitative Free-Response Analysis: An inductive coding approach was used to analyze free-response data. Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words. N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded. Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
3 Results
The undergraduate cohort included 115 respondents across three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other disciplines. The graduate cohort response count was 102 and comprised solely of computer science majors. The full breakdown between sample to response counts can be seen in Table 1.
Figure 2: Response counts by degree and level
Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, seen in Figure 3, with undergraduate computer science students slightly more likely to agree at 21%. Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively. When asked about the limitations of generative AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%. Generative AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
Figure 3: Responses on GenAI awareness.
Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Figure 4 shows undergraduate computer science students have the highest percentage at 82%, followed by computer science graduates at 77%. Other majors report 56%, while engineering students have the lowest percentage at 38%.
Figure 4: Responses on if job skills will change due to GenAI.
Of those that answered “Yes” to job skills changing, there was a follow up question of how soon they believe the changes will occur. Figure 5 shows computer science graduates had the highest expectation of changes occurring the soonest. Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer. The Fall Undergraduate Engineering students mostly expect changes within 5 or 10 years and just 17% expected changes to occur within 3 years.
Figure 6 covers the remaining three questions on GenAI preparedness. Engineering students reported the highest satisfaction with GenAI discussions at 57% and the greatest confidence that professors considered GenAI at 64%. In contrast, Computer Science students had the lowest satisfaction with 38% and the least confidence in faculty engagement at 35%. Despite this, confidence in curriculum preparation was high across all groups, with Engineering leading at 79%. These results suggest stronger GenAI integration in Engineering curricula, while Computer Science students perceive gaps in discussion and faculty consideration.
Figure 5: Respondent counts by degree and cohort.
Figure 6: Responses on preparedness
Concern: The majority view GenAI’s career impact positively, seen in Figure 7, with computer science graduates highest at 81% and undergraduate Engineers, Computer Science, and Other lower with 70%, 64%, and 63% respectively. Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%. Other respondents notably lower at 67%. Even with excitement and optimism about positive career impacts, concern was even greater across all undergraduate degree categories. While less, the computer science graduate cohort still showed concern with 65% concerned with GenAI’s.
Figure 7: Responses on concern
Free Response: In addition to the structured survey responses, a notable portion of participants contributed to the optional open-ended questions. These qualitative insights deepen understanding of the perspectives behind the patterns and are summarized below. 40–50% of undergraduates and graduate students responded to the institutional recommendation prompt, and 23–25% provided input on the additional thoughts prompt. Participation rates for the free response can be seen in Table 3, showing non-Computer Science students having the highest participation. The Fall undergraduate free responses were analyzed as a whole.
Table 3: Optional Free Response Participation by Question & Degree Category
Category Institutional Recommendation Total Additional Thoughts Total Survey Response Total
Undergraduate Other 17 11 38
Undergraduate Engineer 24 11 49
Undergraduate Computer Science 6 5 29
Graduate Computer Science 52 27 102
Figure 8: Topics raised in free response for institutional recommendations
The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use GenAI tools (Figure 8. About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses. Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses. There were fewer who responded to the optional ”additional thoughts” question, those who did showed clear emotional tone(Figure 9). Over half of undergraduates expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to express positivity. Additional thoughts covered a broader range of themes(Figure 10). The computer science graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
Figure 9: Percentage of free responses with either positive or negative emotions conveyed
Figure 10: Topics raised in free response additional thoughts
These open-ended responses provide rich context to the survey data and are further explored in the Discussion. While the first prompt focused on GenAI’s utility, the second invited any remaining reflections from participants.
4 Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers. Although surveys were distributed uniformly to ensure equal access across cohorts, the potential for selection bias and low response frequencies remains a limitation. Response counts were modest compared to sample size, and some subgroups—particularly within non-CS or less tech-aligned majors—were underrepresented. As a technology-focused institution, [Institution Name Redacted for Review]’s student population may also have greater exposure to emerging technologies, including GenAI, than students at other institutions. Even with these constraints, the findings offer important early insight into how GenAI is shaping student experiences, perceptions, and expectations during a critical period of educational and technological transformation.
Awareness: the uncertainty in early adopters
While overall awareness of GenAI was high across all groups, levels of understanding and confidence varied. Computer science undergraduates were the most likely to consider GenAI a short-lived trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%). Despite this skepticism, nearly all respondents across disciplines viewed GenAI as a meaningful advancement, with agreement ranging from 89% to 100%. Experience using GenAI tools also varied, with graduate students reporting the highest exposure (93%) and engineering undergraduates the lowest (78%). Responses regarding the limitations of GenAI’s potential were more common among engineering and graduate students. For the graduate cohort comprising of computer science, this could suggest increased exposure may come with greater critical awareness of its boundaries. For the undergraduate engineering students, skepticism may come from the belief that GenAI’s application is limited specifically for their degree. These nuances are echoed in free responses, where students shared diverse sentiments:
“Because aerospace engineering is pretty specialized and oftentimes requires very difficult/complex work, I am ultimately unsure how GenAI would be used in my field, if at all. However, I trust it can be a valuable tool in the future. I don’t feel like GenAI has any negative impacts on my future.” — Undergraduate Engineer
“I think it has limited application in anything with more advanced logic or understanding.” — Undergraduate Engineer
“I think it has limited potential because GenAI currently cannot distinct whether the information it gives is true or not.” — Undergraduate Engineer
“It’s very uncertain and that’s what makes me slightly afraid. I also don’t understand it very well, which I should work on.” — Undergraduate Computer Science
As seen in some of these responses, there is skepticism of GenAI achieving beyond what the current potential shows, lending to an experience of technology moving faster beyond students’ grasp. Many students rely on self-directed learning, coursework, and personal experiences to form their perspectives on GenAI. This gap in structured AI education leads to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
In the optional free responses, 50% of the students expressed negative emotions, including frustration, fear, and worry, while 20% conveyed uncertainty, with overlapping sentiments highlighting the complex emotional landscape students navigate:
“I work in security. I am not excited about what’s in store.” — Undergraduate Other
“I’m scared it’s going to make my career not what I want it anymore.” — Undergraduate Other
These comments reflect fear and uncertainty about GenAI’s impact, amplified by a lack of classroom discussion. Without institutional guidance, students are left to navigate these changes on their own during critical educational years. Some expressed dissatisfaction with the current state of awareness in their programs overall:
“Even addressing its existence in class would be a step forward.” — Undergraduate Engineer
These emotional responses and findings reveal a misalignment between perceived career impact and readiness and signal a broader psychosocial landscape in flux where students must construct academic identity and career expectations while navigating an unregulated and evolving technology. For a transformative tool like GenAI, students are actively seeking guidance to understand its potential, limits, and ethical implications.
Preparedness: addressing the elephant in the classroom
While students recognize GenAI’s growing influence on their careers, many remain unsure whether their education is preparing them for the changes ahead. Unlike future students who may benefit from GenAI-integrated curricula, the GenAI Generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
This preparedness gap is reflected in how different groups perceive GenAI’s impact on job skills. There is a clear pattern that begins to emerge: those who feel prepared in their degree are most likely to expect a positive career impact.
Where computer science 77% of graduates and 82% undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
Despite these varied views, all cohorts reported relatively high confidence in their curricula—possibly reflecting trust in their degree or Institutions’ reputation. Still, engineering students stand out for their skepticism: they are the least likely to expect GenAI-driven skill changes, report the lowest satisfaction with classroom discussions, and express the least confidence in faculty engagement. This gap signals a need for deeper integration and clearer guidance.
The free response data adds depth to these findings. Despite overall confidence in their curriculum, in responding to the first institutional recommendations response question, over half of students wanted to be taught how to use GenAI. Many also called for stronger emphasis on fundamentals, critical thinking, and ethical considerations:
“I think focusing on the idea of working alongside GenAI on how to take advantage of it to produce better, faster results, but acknowledge its limitations […] and discussing ethical and other ramifications of using AI in certain situations. — Undergraduate Other
“It is a ”tool” that can be used to aid students’ work when used responsibly, and can be integrated in courses like other software tools, such as matlab, cad softwares, etc. but should not replace the fundamentals of engineering that is already being taught, such as hand calculations and understanding the theory.” — Undergraduate Engineer
“Strong implementation in to curriculum and learning tools to learn to work with AI and not treat it as a monster.” — Undergraduate Engineer
“[…] the more adapted we are to it and prepared for its growth/advancement, the better off we’ll be to use it as a tool, rather than as a toy. — Undergraduate Other
While many responses showed a proactive approach to GenAI integration, others expressed skepticism or disappointment:
“[…]I’m disappointed that tech won’t prepare me for the near advent of those tools during my classes.” — Undergraduate Engineer
“As far as I can tell, it’s a fad for almost all possible proposed use cases. Until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
“Show the downsides and flaws of things like gpt-4, discuss environmental impacts, and don’t use it. GenAI is built on stolen data sets and harms artists. Stop endorsing its use. — Undergraduate Other
As GenAI begins to enter technical curricula, students in other disciplines risk being left behind. This gap underscores a fundamental pedagogical challenge: while students recognize GenAI’s relevance, they often do not see it meaningfully integrated into their learning environments, leaving students to independently navigate tools that carry both potential and risk. This lack of structure points to an urgent need for faculty guidance and training, intentional curriculum design, and interdisciplinary GenAI literacy across academic programs. These efforts are essential to prepare a more equitable and capable generation for a GenAI-driven future.
Concern: use at what cost?
Student concerns about GenAI range from ethical implications and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience. The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%). This likely reflects the integration of GenAI into their curriculum and their confidence using such tools.
“[…] GenAI may be able to do the bulk of the work when designing something functional and usable, so the focus can now be on designing technology that people like and that benefits them. I think GenAI can help me do what I like, which would benefit people and society.“ — Graduate Computer Science
“I’m excited for the changes and technological advances it will bring.“ — Graduate Computer Science
In contrast, positive outlook drops to 64% among undergraduate computer science students, reinforcing how exposure shapes perception.
Among undergraduates, belief in a positive career impact ranges from 63% to 70%, with engineering students on the higher end. Excitement is highest among computer science (75%) and engineering students (77%), while the “Other” cohort remains steady at 67%. Even when students are unsure about GenAI’s direct impact on their careers, many still express enthusiasm about its broader potential. Some specific positive takes included:
“I think less people should have to work for as long and more leisure and volunteering time will be helpful for mental health and balance.” — Undergraduate Engineer
“GenAI is the beginning of the 3rd “industrial” revolution. My life has changed a lot from ChatGPT.” — Undergraduate Other
“I think it is important to emphasize both positives and negatives of GenAI whenever mentioned. ” — Undergraduate Computer Science
Free responses reflect strong concern, especially in the the open ended qustion asking for additional thoughts. Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
“I think it will actively damage the quality of work and data produced[…]” — Undergraduate Engineer
“To be honest, I really do not care. I feel like GenAI is going to obliterate most jobs. It almost seems inevitable.[…]The only thing holding me from hopelessness is (blind) faith that my role as a researcher will not be rendered obsolete.” — Undergraduate Computer Science
“I think we are on a precipice of either ruin or greatness. Regardless of how we use it, GenAI will become a powerful tool and have an impact greater than any singular tool since humans gained control of fire.[…] ”— Undergraduate Other
Some students called for regulation:
“[…]we must ensure that there are precautions before the ai reaches a level where its misuse could pose a legitimate harm towards society and overall human advancement.” — Undergraduate Other
Others expressed concern about government use:
“I think the governments will start using genai to produce propoganda in the coming years (if they haven’t already) to manipulate public opinion. Or maybe just politicians with bot accounts blindly supporting them and arguing with dissenters on social media.” — Undergraduate Engineer
And some of the more succinct, poignant responses:
“Goodbye middle class, we doomed.” — Undergraduate Computer Science
A compelling insight lies in who chose to share their perspectives in the optional free-response section. Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduates. This could suggest that students who may not encounter GenAI in their formal coursework still feel strongly about its impact and relevance.
Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement. Students are not only grappling with how to use GenAI but also with what kind of future it is shaping.
5 Conclusion
This study contributes to the growing discourse on generative AI in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance. It introduces the term GenAI Generation to describe those in their formative educational years during the rise of GenAI, students whose academic development is unfolding alongside the rapid adoption of these tools. Through analysis of over 250 survey responses, the findings reveal not only patterns of perceived readiness and ethical concern, but also a broader psychosocial landscape and pedagogical gap. Students are engaging with GenAI amid unclear norms, forming academic identities and career expectations as institutions lag behind. These insights offer timely direction for educators and policymakers seeking to design inclusive, developmentally aligned, and ethically grounded approaches to GenAI integration.
The GenAI Generation faces a rapidly evolving landscape shaped by both excitement about GenAI’s potential and concern about preparedness, ethics, and long-term implications. This study reveals a growing divide in student confidence and readiness, shaped more by exposure and self-directed integration of GenAI than by formal curriculum. However, structured incorporation is inconsistent across disciplines, highlighting a broader institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift. The divide is no longer simply between STEM and non-STEM fields, but between those with and without access to meaningful engagement with emerging technologies.
The GenAI Generation is distinct in its developmental context. These students are not only adapting to powerful new tools, but doing so without established norms or guidance. As early adopters, they must balance traditional skills with new competencies while grappling with ethical dilemmas, shifting workforce expectations, and unknown cognitive impacts. Their adaptability is being tested in real time.
Students want more than access, they are asking for structure. They seek guidance on using GenAI responsibly, developing critical thinking around AI-generated outputs, and building a strong foundation in ethical reasoning. Their feedback calls for a holistic approach that prepares them not only to use GenAI, but to question and shape it.
As the first cohort to learn and grow in a GenAI-centric world, the experiences of the GenAI Generation will set the tone for future educational trajectories. Their voices point to an urgent need for GenAI literacy that spans disciplines and prepares both students and faculty to engage with these tools thoughtfully and equitably. They envision a future where GenAI is not a source of confusion or inequity, but a tool for growth—integrated into education in ways that are ethical, empowering, and human-centered.
6 List of abbreviations
Abbreviation Full Form
GenAI Generative Artificial Intelligence
AI Artificial Intelligence
IRB Institutional Review Board
OMSCS Online Master of Science in Computer Science
MSCS Master of Science in Computer Science
OAE [Institution Name Redacted for Review] Office of Academic Effectiveness
STEM Science, Technology, Engineering, and Mathematics
CS Computer Science (used in tables/figures for grouping)
7 Future Considerations
To better understand the evolving impact of GenAI on student development and educational outcomes, future research should consider longitudinal studies that track student perceptions, readiness, and concerns over time. As GenAI tools rapidly advance, students’ experiences and attitudes will likely shift in response to both technological innovation and institutional adaptation. Capturing these changes across semesters or academic years would provide valuable insight into how exposure, policy, and pedagogy shape preparedness and perception in the long term.
In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on GenAI tools, such as ChatGPT, for learning and problem-solving. Cognitive offloading may free students to focus on higher-order thinking, but it also raises questions about memory retention, deep learning, and critical reasoning. Understanding how sustained use of GenAI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support—not replace—core learning processes.
There is also a need for targeted dissemination and outreach to disciplines beyond computer science. This study found that students in non-computer science fields, particularly those in the ”Other” category, demonstrated strong emotional responses and a desire for inclusion in GenAI discourse, despite limited curricular exposure. Expanding GenAI literacy and integration into these areas can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
8 Declarations
8.1 Funding
This study was conducted without external funding. No funding agency was involved in the design, data collection, analysis, or publication of this research.
8.2 Funding Statement
8.3 Consent to Publish declaration
Not applicable
8.4 Consent to Participate
This research study was Institutional Review Board (IRB) approved. Respondents provided informed consent.
8.5 Ethics declaration
This research study was Institutional Review Board (IRB) approved. The study collected students’ perceptions of their future career preparedness in light of the potential impact of generative artificial intelligence (GenAI) through a survey. The survey only collected responses for students located within the United States, and 18 years of age or older as per IRB instruction. Respondents provided informed consent. Survey questions were designed to protect participant identity. Participation was completely voluntary, and had the ability to opt out at any time. No benefit or compensation was provided. We complied with any applicable laws and regulations regarding confidentiality. The IRB and Office of Human Research Protections reserved the right to look at study records at any time. Contacts to the author and Office of Research Integrity Assurance were provided.